# Exact inference via quasi-conjugacy in two-parameter Poisson–Dirichlet hidden Markov models

### v1

@@ A.4.2 Offline estimation: smoothing and data forecasting
$\Pr(\Pi^{N}=\pi\mid X(t_{N-1})=x)=\mathrm{PSF}_{\alpha,\theta}(\pi)\sum_{\lambda\in L(\pi)}h_{\lambda}(\Delta_{N})\,\frac{P(\lambda\mid x)}{\mathrm{PSF}_{\alpha,\theta}(\lambda)},\qquad\pi\in\mathcal{P}_{n_{N}},$ (48)
@@
$\Pr(\Pi^{k}=\pi^{k},\dots,\Pi^{N}=\pi^{N}\mid X(t_{k-1})=x)$ $=\prod_{j=k}^{N}g(\pi^{j}\mid\pi^{j+1:N})\,\sum_{\omega\in\Omega^{k:N}}h_{\omega}(\Delta_{k})\,\frac{P(\omega\mid x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (49)
@@
$g(\pi^{k}\mid\pi^{k+1:N})=\sum_{\omega\in\Omega^{k+1:N}}h_{\omega}(\Delta_{k+1})\,\mathrm{CRP}_{\omega\uparrow\pi^{k}}.$ (51)
@@
$h_{\omega}(\Delta_{k})=\sum_{\mu\in\mathrm{coag}(\Omega^{k+1:N},\pi^{k})}\left[\sum_{\omega^{\prime}\in\Omega^{k+1:N}}\frac{h_{\omega^{\prime}}(\Delta_{k+1})\,\mathrm{CRP}_{\omega^{\prime}\uparrow\pi^{k}}}{g(\pi^{k}\mid\pi^{k+1:N})}\,\Pr(\mu\mid\omega^{\prime},\pi^{k})\right]p^{\downarrow}_{\mu,\omega}(\Delta_{k}),$ (52)
with $\mathrm{CRP}_{\omega^{\prime}\uparrow\pi^{k}}$ as in (33) and $\Pr(\mu\mid\omega^{\prime},\pi^{k})$ as in (34).
@@
For $k=0,\dots,N$, the smoothed law of $X(t_{k})\mid\Pi^{0:N}$ belongs to $\mathcal{F}$ and can be written as
$X(t_{k})\mid\lambda(t)\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t)},\qquad\lambda(t)\mid\Pi^{0:N}\sim\sum_{(\lambda,\omega)\in\Lambda_{0:k}\times\Omega^{k:N}}w_{\lambda,\omega}\,\mathrm{Coag}_{\alpha,\theta}(\lambda,\omega),$ (54)
@@
The same structure yields $\Pr(X(t_{k}+\delta)\in\cdot\mid\Pi^{0:N})$ for $0<\delta<\Delta_{k+1}$.
@@
For $t_{k}+\delta\in(t_{k},t_{k+1})$, $\Pr(X(t_{k}+\delta)\in{\rm d}x\mid\Pi^{0:N})\in\mathcal{F}$ and
$\Pr(X(t_{k}+\delta)\in{\rm d}x\mid\Pi^{0:N})=\sum_{\mu\in\mathrm{coag}\big(L(\Lambda_{0:k}),\ \Omega^{k:N}\big)}s_{\mu}(\delta,\Delta_{k+1})\,\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x),$ (56)
@@
$s_{\mu}(\delta,\Delta_{k+1})\ \propto\ \sum_{\eta\in L(\Lambda_{0:k})}\sum_{\omega\in\Omega^{k:N}}w_{\eta,\omega}(\delta,\Delta_{k+1})\,\Pr(\mu\mid\eta,\omega),$ (57)
@@
We can also simulate future (or interpolated) partitions $\Pi^{\prime}(t)$ of size $n$ conditional on $\Pi^{0:N}$.
@@
$\Pi^{\prime}(t)\mid\lambda(t)\ \sim\ \mathrm{CRP}_{\lambda(t)\uparrow\bullet},\qquad\lambda(t)\ \sim\ \{w_{\lambda}\},$ (59)
where $\{w_{\lambda}\}$ are the mixture weights for $\Pr(X(t)\in{\rm d}x\mid\Pi^{0:N})$, obtained from: Theorem 2 if $t>t_{N}$; Theorem 3 if $t=t_{N}$; Theorem 4 if $t=t_{k}<t_{N}$; or Corollary 2 if $t\in(t_{k},t_{k+1})$.
Algorithm 8 then samples Young diagrams conditionally on $\pi^{0:N}$. Equation (59) can be seen as a temporally informed $\mathrm{CRP}$, where temporal dependence is encoded by the weights $\{w_{\lambda}\}$; the overall clustering structure is unknown within the finite set of plausible coagulations.
A.4.3 Proofs
Proof of Proposition 4. We give here a direct probabilistic proof which highlights the role of the random sampling. See also feng2011functional; feng2010poisson for previous discussion of similar results. By (24), we can write
$P(\omega|x)P(\gamma|x)=\mathrm{Pr}_{x}(\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma)$
which factorizes since the two samples are conditionally independent given $x$. By construction, $\mathrm{Pr}_{x}(\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma|\Pi_{1:n+m}=\mu)$ is in fact independent of $x$ and coincides with the sum of joint hypergeometric probabilities in (4). Then
$\mathrm{Pr}_{x}($ $\,\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma)$ (60)
@@ A.4.3 Proofs
Proof of Proposition 5. The law of a partition drawn from a two-parameter Chinese restaurant process can be written by integrating out the $\mathrm{PD}_{\alpha,\theta}$-distributed random measure, yielding the de Finetti representation
$\Pr(\Pi_{n+1:n+m}=\gamma\mid\Pi_{1:n}=\omega)=$ $\,\int_{\nabla}P(\gamma|x)\,\mathrm{PD}_{\alpha,\theta}^{\omega}({\rm d}x),$ (64)
@@
Using (64), the left-hand side of Proposition 5 can be written as a mixture,
@@
where $\mathrm{Pr}_{x}$ denotes the law of random partitions induced by subsequent draws from $x\in\nabla$; cf. (24). Using (64), the previous equals
@@
since $P(\omega|x)=\mathrm{Pr}_{x}(\Pi_{1:n}=\omega)$. Now, conditionally on $\Pi_{1:n+m}$ the subpartitions $\Pi_{1:n}$ and $\Pi_{n+1:n+m}$ are independent of $x$, (cf. proof of Proposition 4) and their conditional joint law is the hypergeometric in Proposition 4. Hence we have
@@
$=\mathrm{Pr}_{x}(\Pi_{1:n+m}=\mu)\mathcal{H}(\omega,\gamma|\mu),$
@@
$\frac{\mathcal{H}(\omega,\gamma|\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\int_{\nabla}\mathrm{Pr}_{x}(\Pi_{1:n+m}=\mu)\mathrm{PD}_{\alpha,\theta}({\rm d}x),$
and the claim is now implied by Kingman’s representation theorem (kingman_representation_1978; pitman1995exchangeable).∎
@@
Due to the duality between the $\mathrm{PD}_{\alpha,\theta}$ diffusion and the death process described in Section A.3, (67) becomes
@@
and since $D_{t}$ is a death process on $\mathcal{P}$, we obtain
@@
$|\varphi_{t}(\nu_{N|0:N})(A)$ $-\mathrm{PD}_{\alpha,\theta}(A)|\leq\sum_{\eta\in L(\Lambda)}v_{\eta}(t)\bigg|\mathrm{PD}_{\alpha,\theta}^{\omega}(A)-\mathrm{PD}_{\alpha,\theta}(A)\bigg|$
@@
$\,+\sum_{\omega\in L(\Lambda_{0:N}):|\omega|>1}v_{\omega}(t)\bigg|\mathrm{PD}_{\alpha,\theta}^{\omega}(A)-\mathrm{PD}_{\alpha,\theta}(A)\bigg|.$
@@
$v_{\omega}(t)=\sum_{\lambda\in\Lambda_{0:N}}w_{\lambda}\mathcal{H}(\omega|\lambda)d_{|\lambda|,|\omega|}^{\theta}(t)$
the fact that $d_{|\lambda|,|\omega|}^{\theta}(t)$ is the transition probability of a death-process in $\mathbb{Z}_{+}$ with absorption in $(1)$ implies that $v_{\omega}(t)\rightarrow 0$ as $t\rightarrow\infty$ for all $\omega$ such that $|\omega|>1$. The result is now implied by the fact that $\mathrm{PD}_{\alpha,\theta}^{(1)}\overset{d}{=}\mathrm{PD}_{\alpha,\theta}$ (see Remark 1).∎
@@
$\sum_{\omega\in L(\Lambda_{0:k-1})}$ $\,v_{\omega}\sum_{\mu\in\mathrm{coag}(\omega,\pi^{k})}\Pr(\mu,\pi^{k}|\omega)=\sum_{\omega\in L(\Lambda_{0:k-1})}v_{\omega}\mathrm{CRP}_{\omega\uparrow\pi^{k}}$
and multiply and divide by $\mathrm{CRP}_{\omega\uparrow\pi^{k}}$ to obtain
$\sum_{\omega\in L(\Lambda_{0:k-1})}\frac{v_{\omega}\mathrm{CRP}_{\omega\uparrow\pi^{k}}}{\sum_{\omega^{\prime}}v_{\omega^{\prime}}\mathrm{CRP}_{\omega^{\prime}\uparrow\pi^{k}}}\sum_{\mu\in\mathrm{coag}(\omega,\pi^{k})}\Pr(\mu|\omega,\pi^{k})\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x).$
@@
Proof of Proposition 6. Let us avoid to write the random partitions $\Pi^{k:N}$ in $\mathrm{Pr}_{\alpha,\theta}$ to ease the notation. We proceed by induction on $k$. First note that
@@
where in the last identity we have used the duality (recall Section A.3). Now, since
$\mathrm{Pr}_{\alpha,\theta}(\pi^{k:N}|X(t_{k-1})=x)$ $=\int_{\nabla}\mathrm{Pr}_{\alpha,\theta}(\pi^{k:N}|X(t_{k})=z)p_{\Delta_{k}}(x,{\rm d}z)$ (75)
$=\int_{\nabla}P(\pi^{k}|z)\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k})=z)p_{\Delta_{k}}(x,{\rm d}z),$ (76)
@@
$\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k})=x)=\prod_{j=k+1}^{N}g(\pi^{j}|\pi^{j+1:N})\sum_{\omega\in\Omega^{k+1:N}}h_{\omega}(\Delta_{k+1})\frac{P(\omega|x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (77)
@@
and using the duality (recall Section A.3), we find
@@
$h_{\eta}(\Delta_{k}):=\sum_{\mu\in\mathrm{coag}(\pi^{k},\Omega^{k+1:N})}\left[\sum_{\omega\in\Omega^{k+1:N}}\frac{h_{\omega}(\Delta_{k+1})\mathrm{CRP}_{\omega\uparrow\pi^{k}}}{g(\pi^{k}|\pi^{k+1:N})}\Pr(\mu|\omega,\pi^{k})\right]p^{\downarrow}_{\mu,\eta}(\Delta_{k}),$ (79)
@@
$\mathrm{Pr}_{\alpha,\theta}(\pi^{k:N}|X(t_{k-1})=x)=\prod_{j=k}^{N}g(\pi^{j}|\pi^{j+1:N})\sum_{\omega\in\Omega^{k:N}}h_{\omega}(\Delta_{k})\frac{P(\omega|x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (80)
@@
$\nu_{k|0:N}(A)=$ $\,\mathrm{Pr}_{\alpha,\theta}(X(t_{k})\in A|\pi^{0:N})$ (81)
$\propto$ $\,\int_{A}\mathrm{Pr}_{\alpha,\theta}(X(t_{k})\in{\rm d}x|\pi^{0:k})\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k})=x)$
$=$ $\,\int_{A}\nu_{k|0:k}({\rm d}x)\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k})=x).$
where the random partitions $\Pi^{0:N}$ are not included in $\mathrm{Pr}_{\alpha,\theta}$ just to ease the notation. Hence, up to proportionality, $\nu_{k|0:N}$ factorizes into $\nu_{k|0:k}$ and the likelihood of partitions collected after time $t_{k}$, the latter given in Proposition 6. As the first factor in (81) is the filtering distribution $\nu_{k|0:k}$ in the statement, we now have
@@
$\sum_{\mu\in\mathrm{coag}(\Lambda^{0:k},\ \Omega^{k:N})}\left[\sum_{\begin{subarray}{c}\omega\in\Omega^{k:N}\\ \lambda\in\Lambda_{0:k}\end{subarray}}\frac{w_{\lambda}\mathrm{CRP}_{\lambda\uparrow\omega}h_{\omega}(\Delta_{k+1})}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\Pr(\mu|\lambda,\omega)\right]\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x),$ (87)
which concludes the proof upon noting that $\Pr(\lambda,\omega)=\mathrm{PSF}_{\alpha,\theta}(\lambda)\mathrm{CRP}_{\lambda\uparrow\omega}$, which follows from (64). ∎
@@
$\mathrm{Pr}_{\alpha,\theta}(X(t_{k}+\delta)\in{\rm d}x|\pi^{0:N})$ $\propto\mathrm{Pr}_{\alpha,\theta}(X(t_{k}+\delta)\in{\rm d}x|\pi^{0:k})\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k}+\delta)=x)$
$=\varphi_{\delta}(\nu_{0:k})({\rm d}x)\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k}+\delta)=x).$
where the random partitions $\Pi^{0:N}$ are omitted in $\mathrm{Pr}_{\alpha,\theta}$ just to ease the notation. As already shown in Theorem 2, the propagation of $\nu_{k|0:k}$ is
@@
In order to find the likelihood $\mathrm{Pr}_{\alpha,\theta}(\pi^{k+1:N}|X(t_{k}+\delta)=x)$ one can proceed as in the proof of Proposition 6, but replacing $p^{\downarrow}_{\mu,\eta}(\Delta_{k+1})$ with $p^{\downarrow}_{\mu,\eta}(\Delta_{k+1}-\delta)$.∎
@@
$\int_{\nabla}P(\pi|x)\mathrm{Pr}_{\alpha,\theta}(X(t)\in{\rm d}x|\pi^{0:N})=\sum_{\lambda\in\Lambda_{t}}w_{\lambda}\int_{\nabla}P(\pi|x)\mathrm{PD}_{\alpha,\theta}^{\lambda}({\rm d}x)$

### v2

@@ A.3.2 Offline estimation: smoothing and data forecasting
$\Pr(\Pi^{N}=\pi|X(t_{N-1})=x)=\mathrm{PSF}_{\alpha,\theta}(\pi)\sum_{\lambda\in L(\pi)}h_{\lambda}(\Delta_{N})\,\frac{P(\lambda|x)}{\mathrm{PSF}_{\alpha,\theta}(\lambda)},\qquad\pi\in\mathcal{P}_{n_{N}},$ (51)
@@
$\Pr(\Pi^{k}=\pi^{k},\dots,\Pi^{N}=\pi^{N}|X(t_{k-1})=x)$ $=\prod_{j=k}^{N}g(\pi^{j}|\pi^{j+1:N})\,\sum_{\omega\in\Omega^{k:N}}h_{\omega}(\Delta_{k})\,\frac{P(\omega|x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (52)
@@
$g(\pi^{k}|\pi^{k+1:N})=\sum_{\omega\in\Omega^{k+1:N}}h_{\omega}(\Delta_{k+1})\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k}).$ (54)
@@
$h_{\omega}(\Delta_{k})=\sum_{\mu\in\mathrm{coag}(\Omega^{k+1:N},\pi^{k})}\left[\sum_{\omega^{\prime}\in\Omega^{k+1:N}}\frac{h_{\omega^{\prime}}(\Delta_{k+1})\mathrm{PSF}_{\alpha,\theta}^{\omega^{\prime}}(\pi^{k})}{g(\pi^{k}|\pi^{k+1:N})}\,\Pr(\mu|\omega^{\prime},\pi^{k})\right]p^{\downarrow}_{\mu,\omega}(\Delta_{k}),$ (55)
with $\mathrm{PSF}_{\alpha,\theta}^{\omega^{\prime}}$ as in (36) and $\Pr(\mu|\omega^{\prime},\pi^{k})$ as in (37).
@@
For $k=0,\dots,N$, the smoothed law of $X(t_{k})|\Pi^{0:N}$ belongs to $\mathcal{F}$ and can be written as
$X(t_{k})|\lambda(t)\sim\mathrm{PD}_{\alpha,\theta}^{\lambda(t)},\qquad\lambda(t)|\Pi^{0:N}\sim\sum_{(\lambda,\omega)\in\Lambda_{0:k}\times\Omega^{k:N}}w_{\lambda,\omega}\,\mathrm{Coag}_{\alpha,\theta}(\lambda,\omega),$ (57)
@@
The same structure yields $\Pr(X(t_{k}+\delta)\in\cdot|\Pi^{0:N})$ for $0<\delta<\Delta_{k+1}$.
@@
For $t_{k}+\delta\in(t_{k},t_{k+1})$, $\Pr(X(t_{k}+\delta)\in{\rm d}x|\Pi^{0:N})\in\mathcal{F}$ and
$\Pr(X(t_{k}+\delta)\in{\rm d}x|\Pi^{0:N})=\sum_{\mu\in\mathrm{coag}\big(L(\Lambda_{0:k}),\ \Omega^{k:N}\big)}s_{\mu}(\delta,\Delta_{k+1})\,\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x),$ (59)
@@
$s_{\mu}(\delta,\Delta_{k+1})\ \propto\ \sum_{\eta\in L(\Lambda_{0:k})}\sum_{\omega\in\Omega^{k:N}}w_{\eta,\omega}(\delta,\Delta_{k+1})\,\Pr(\mu|\eta,\omega),$ (60)
@@
We can also simulate future (or interpolated) partitions $\Pi^{\prime}(t)$ conditional on $\Pi^{0:N}$.
@@
$\Pi^{\prime}(t)|\lambda(t)\ \sim\ \mathrm{PSF}_{\alpha,\theta}^{\lambda(t)},\qquad\lambda(t)\ \sim\ \{w_{\lambda};\lambda\in\Lambda_{t}\},$ (62)
where $\{w_{\lambda};\lambda\in\Lambda_{t}\}$ are the mixture weights for $\Pr(X(t)\in{\rm d}x|\Pi^{0:N})$, obtained from: Theorem 2 if $t>t_{N}$; Theorem 3 if $t=t_{N}$; Theorem 4 if $t=t_{k}<t_{N}$; or Corollary 2 if $t\in(t_{k},t_{k+1})$.
Algorithm 8 then samples partitions conditionally on $\pi^{0:N}$. As in Section A.2.3, Equation (62) can be seen as a temporally informed Chinese Restaurant Process, where temporal dependence is encoded by the weights $\{w_{\lambda};\lambda\in\Lambda_{t}\}$; the overall clustering structure is unknown within the finite set of plausible coagulations
A.3.3 Proofs
Proof of Proposition 4. We give here a direct probabilistic proof which highlights the role of the random sampling. See also feng2011functional; feng2010poisson for previous discussion of similar results. Conditionally on $x\in\nabla$, $\Pi_{1:n}$ and $\Pi_{n+1:n+m}$ are respectively the clusterings of the first $n$ and of the next $m$ samples of Kingman’s paintbox process from $x$. Then by definiton of the paintbox process the joint probability factorizes as
$\mathrm{Pr}_{x}(\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma)=\mathrm{Pr}_{x}(\Pi_{1:n}=\omega)\mathrm{Pr}_{x}(\Pi_{n+1:n+m}=\gamma)=P(\omega|x)P(\gamma|x).$
Moreover, it is plain to see that by definition of the paintbox process, conditionally on the partition $\Pi_{1:n+m}$ obtained by all $n+m$ samples, the probability
$\mathrm{Pr}_{x}(\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma|\Pi_{1:n+m}=\mu)$
is in fact independent of $x$ and coincides with the sum of joint hypergeometric probabilities in (4). Then
$P(\omega|x)P(\gamma|x)$ $=\mathrm{Pr}_{x}(\Pi_{1:n}=\omega,\Pi_{n+1:n+m}=\gamma)$ (63)
@@ A.3.3 Proofs
Proof of Proposition 5. By virtue of Kingman’s representation 29, we known that the marginal law on exchangeable partitions can be represented as a mixture of paintboxes. In particular, as in 31, the Chinese Restaurant Process is obtained by integrating out $X\sim\mathrm{PD}_{\alpha,\theta}$ from the law of the paintbox of $X$. Then
$\Pr(\Pi_{n+1:n+m}=\gamma|\Pi_{1:n}=\omega)=$ $\,\int_{\nabla}P(\gamma|x)\,\mathrm{PD}_{\alpha,\theta}^{\omega}({\rm d}x),$ (67)
@@
See also griffiths2024dual. Using (67), the left-hand side of Proposition 5 can be written as a mixture,
@@
where $\mathrm{Pr}_{x}$ denotes the law of the paintbox from $x\in\nabla$; cf. (25). Using (67), the previous equals
@@
since by definition $P(\omega|x)=\mathrm{Pr}_{x}(\Pi_{1:n}=\omega)$. Now, conditionally on $\Pi_{1:n+m}$ the subpartitions $\Pi_{1:n}$ and $\Pi_{n+1:n+m}$ are independent of $x$, (cf. proof of Proposition 4) and their conditional joint law is the hypergeometric in Proposition 4. Hence we have
@@
$=P(\mu|x)\mathcal{H}(\omega,\gamma|\mu),$
@@
$\frac{\mathcal{H}(\omega,\gamma|\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\int_{\nabla}P(\mu|x)\mathrm{PD}_{\alpha,\theta}({\rm d}x)=\mathcal{H}(\omega,\gamma|\mu)\frac{\mathrm{PSF}_{\alpha,\theta}(\mu)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$
which follows once again from 31.∎
@@
Due to the duality between the $\mathrm{PD}_{\alpha,\theta}$ diffusion and the death process described in Section A.2.3, (70) becomes
@@
and since $D(t)$ is a death process on $\mathcal{P}$, we obtain
@@
$|\varphi_{t}(\nu_{N|0:N})(A)$ $-\mathrm{PD}_{\alpha,\theta}(A)|\leq\sum_{\eta\in L(\Lambda)}v_{\eta}(t)\bigg|\mathrm{PD}_{\alpha,\theta}^{\eta}(A)-\mathrm{PD}_{\alpha,\theta}(A)\bigg|$
@@
$\,+\sum_{\eta\in L(\Lambda_{0:N}):|\eta|>1}v_{\eta}(t)\bigg|\mathrm{PD}_{\alpha,\theta}^{\eta}(A)-\mathrm{PD}_{\alpha,\theta}(A)\bigg|.$
@@
$v_{\eta}(t)=\sum_{\lambda\in\Lambda_{0:N}}w_{\lambda}\mathcal{H}(\eta|\lambda)d_{|\lambda|,|\eta|}^{\theta}(t)$
the fact that $d_{|\lambda|,|\eta|}^{\theta}(t)$ is the transition probability of a death-process in $\mathbb{Z}_{+}$ with absorption in $(1)$ implies that $v_{\eta}(t)\rightarrow 0$ as $t\rightarrow\infty$ for all $\eta$ such that $|\eta|>1$. The result is now implied by the fact that $\mathrm{PD}_{\alpha,\theta}^{(1)}\overset{d}{=}\mathrm{PD}_{\alpha,\theta}$ (see Remark 1).∎
@@
$\sum_{\omega\in L(\Lambda_{0:k-1})}$ $\,v_{\omega}\sum_{\mu\in\mathrm{coag}(\omega,\pi^{k})}\Pr(\mu,\pi^{k}|\omega)=\sum_{\omega\in L(\Lambda_{0:k-1})}v_{\omega}\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})$
and multiply and divide by $\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})$ to obtain
$\sum_{\omega\in L(\Lambda_{0:k-1})}\frac{v_{\omega}\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})}{\sum_{\omega^{\prime}}v_{\omega^{\prime}}\mathrm{PSF}_{\alpha,\theta}^{\omega^{\prime}}(\pi^{k})}\sum_{\mu\in\mathrm{coag}(\omega,\pi^{k})}\Pr(\mu|\omega,\pi^{k})\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x).$
@@
Proof of Proposition 6. Let us avoid to write the random partitions $\Pi^{k:N}$ in $\Pr$ to ease the notation. We proceed by induction on $k$. First note that
@@
where in the last identity we have used the duality (recall Section A.2.3). Now, since
$\Pr(\pi^{k:N}|X(t_{k-1})=x)$ $=\int_{\nabla}\Pr(\pi^{k:N}|X(t_{k})=z)p_{\Delta_{k}}(x,{\rm d}z)$ (78)
$=\int_{\nabla}P(\pi^{k}|z)\Pr(\pi^{k+1:N}|X(t_{k})=z)p_{\Delta_{k}}(x,{\rm d}z),$ (79)
@@
$\Pr(\pi^{k+1:N}|X(t_{k})=x)=\prod_{j=k+1}^{N}g(\pi^{j}|\pi^{j+1:N})\sum_{\omega\in\Omega^{k+1:N}}h_{\omega}(\Delta_{k+1})\frac{P(\omega|x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)},$ (80)
@@
and using the duality (recall Section A.2.3), we find
@@
$h_{\eta}(\Delta_{k}):=\sum_{\mu\in\mathrm{coag}(\pi^{k},\Omega^{k+1:N})}\left[\sum_{\omega\in\Omega^{k+1:N}}\frac{h_{\omega}(\Delta_{k+1})\mathrm{PSF}_{\alpha,\theta}^{\omega}(\pi^{k})}{g(\pi^{k}|\pi^{k+1:N})}\Pr(\mu|\omega,\pi^{k})\right]p^{\downarrow}_{\mu,\eta}(\Delta_{k}),$ (82)
@@
$\Pr(\pi^{k:N}|X(t_{k-1})=x)=\prod_{j=k}^{N}g(\pi^{j}|\pi^{j+1:N})\sum_{\omega\in\Omega^{k:N}}h_{\omega}(\Delta_{k})\frac{P(\omega|x)}{\mathrm{PSF}_{\alpha,\theta}(\omega)}.$ (83)
@@
$\nu_{k|0:N}(A)=$ $\,\Pr(X(t_{k})\in A|\pi^{0:N})$ (84)
$\propto$ $\,\int_{A}\Pr(X(t_{k})\in{\rm d}x|\pi^{0:k})\Pr(\pi^{k+1:N}|X(t_{k})=x)$
$=$ $\,\int_{A}\nu_{k|0:k}({\rm d}x)\Pr(\pi^{k+1:N}|X(t_{k})=x).$
where the random partitions $\Pi^{0:N}$ are not included in $\Pr_{\alpha,\theta}$ just to ease the notation. Hence, up to proportionality, $\nu_{k|0:N}$ factorizes into $\nu_{k|0:k}$ and the likelihood of partitions collected after time $t_{k}$, the latter given in Proposition 6. As the first factor in (84) is the filtering distribution $\nu_{k|0:k}$ in the statement, we now have
@@
$\sum_{\mu\in\mathrm{coag}(\Lambda^{0:k},\ \Omega^{k:N})}\left[\sum_{\begin{subarray}{c}\omega\in\Omega^{k:N}\\ \lambda\in\Lambda_{0:k}\end{subarray}}\frac{w_{\lambda}\mathrm{PSF}_{\alpha,\theta}^{\lambda}(\omega)h_{\omega}(\Delta_{k+1})}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\Pr(\mu|\lambda,\omega)\right]\mathrm{PD}_{\alpha,\theta}^{\mu}({\rm d}x),$ (90)
which concludes the proof upon noting that
$\Pr(\lambda,\omega)=\mathrm{PSF}_{\alpha,\theta}(\lambda)\mathrm{PSF}_{\alpha,\theta}^{\lambda}(\omega)=\mathrm{PSF}_{\alpha,\theta}(\omega)\mathrm{PSF}_{\alpha,\theta}^{\omega}(\lambda),$
which follows from (67). ∎
@@
$\Pr(X(t_{k}+\delta)\in{\rm d}x|\pi^{0:N})$ $\propto\Pr(X(t_{k}+\delta)\in{\rm d}x|\pi^{0:k})\Pr(\pi^{k+1:N}|X(t_{k}+\delta)=x)$
$=\varphi_{\delta}(\nu_{0:k})({\rm d}x)\Pr(\pi^{k+1:N}|X(t_{k}+\delta)=x).$
where the random partitions $\Pi^{0:N}$ are omitted in $\Pr$ just to ease the notation. As already shown in Theorem 2, the propagation of $\nu_{k|0:k}$ is
@@
In order to find the likelihood $\Pr(\pi^{k+1:N}|X(t_{k}+\delta)=x)$ one can proceed as in the proof of Proposition 6, but replacing $p^{\downarrow}_{\mu,\eta}(\Delta_{k+1})$ with $p^{\downarrow}_{\mu,\eta}(\Delta_{k+1}-\delta)$.∎
@@
$\int_{\nabla}P(\pi|x)\Pr(X(t)\in{\rm d}x|\pi^{0:N})=\sum_{\lambda\in\Lambda_{t}}w_{\lambda}\int_{\nabla}P(\pi|x)\mathrm{PD}_{\alpha,\theta}^{\lambda}({\rm d}x)$